# Towards Physical Understanding in Video Generation: A 3D Point Regularization Approach

### v1

@@ Abstract
We present a novel video generation framework that integrates 3-dimensional geometry and dynamic awareness. To achieve this, we augment 2D videos with 3D point trajectories and align them in pixel space. The resulting 3D-aware video dataset, PointVid, is then used to fine-tune a latent diffusion model, enabling it to track 2D objects with 3D Cartesian coordinates. Building on this, we regularize the shape and motion of objects in the video to eliminate undesired artifacts, e.g., nonphysical deformation. Consequently, we enhance the quality of generated RGB videos and alleviate common issues like object morphing, which are prevalent in current video models due to a lack of shape awareness. With our 3D augmentation and regularization, our model is capable of handling contact-rich scenarios such as task-oriented videos. These videos involve complex interactions of solids, where 3D information is essential for perceiving deformation and contact. Furthermore, our model improves the overall quality of video generation by promoting the 3D consistency of moving objects and reducing abrupt changes in shape and motion.
Figure 1: Comparison on Task-Oriented Videos. We present videos generated by different baselines (SVD [6], I2VGen-XL [41], and DynamiCrafter [37]) and compare them with our method. We use the same input conditions for all methods (except that SVD is conditioned only on the image). It can be observed that existing baselines often exhibit severe distortions of hands or objects during human hand-object interactions. In contrast, our method preserves the shapes of both the hand and object during such interactions and ensures smooth transitions throughout the video.
@@ 1 Introduction
The recent development of video diffusion models has attracted significant research interests and seen tremendous progress [6, 32, 14]. With vast amounts of training data and sophisticated architectures, video models have greatly improved in expressiveness and aesthetics [7, 33]. Today, state-of-the-art video models can realistically represent content from the real world with high creativity, enabling various applications in the entertainment industry and scientific research [25]. However, existing video diffusion models primarily focus on improving the appearance of content and the smoothness of motion. The models are trained in a way to understand the movements of 2-dimensional pixels, where the object movements in 3D space is approximated through the changing of RGB values.
@@
As a result, the underlying 3-dimensional shapes and motions, which represent how objects truly exist in the physical world, are not learned well in the existing video generation models. Even worse, we observe that when the video diffusion models neglect the underlying 3D information, they struggle to understand how objects occupy space, change shape and location, and interact with each other, which is essential for generating videos with complex motion. For instance, imaging a situation when perceiving a scene with a human performing specific tasks, e.g., hands tying shoelaces, viewpoint projection and occlusion make it impossible to capture the true 3D shape of the hand and the shoelaces using 2D pixels, complicating the modeling of how they make contact. Without understanding the underlying geometry, objects in the video may change shape arbitrarily or even appear and disappear suddenly, leading to the common issue of object morphing (as example results from prior arts in Fig. 1). This problem is exacerbated for complex objects like humans; without 3D awareness, limbs and body parts may move unnaturally or change shape abruptly, resulting in dislocated arms or twisted bodies.
@@
To address the above issue, we aim to improve video diffusion models towards the better understanding of physical world, thus generating more reasonable shape and motion. Previous works have attempted to control shape and motion directly in 2D space [31, 23]. Nonetheless, without a comprehensive understanding of 3D geometry, these methods are limited to directing pixel movement confined within the 2D plane, making them unable to handle out-of-plane motions or fully resolve 3D dynamics [23]. One straightforward approach could be to model the complete 3D geometry of objects using representations such as NeRF [26] or 3D Gaussian-Splatting [18], essentially leading to 4D generation [3, 43]. While these methods provide accurate 3D understanding, they are computationally intensive and generally limited to generating simple scenes with only a few specific types of objects.
Figure 2: Training Pipeline Overview. We sample video-point pairs, concatenate them in channel dimensions and used to train a UNet. In addition to standard condition and latent cross attention, we further add cross attention between video and point in corresponding channels for a better alignment between the two modalities. Furthermore, the 3D information from the points is utilized to regularize the RGB video generation by applying a misalignment penalty to the video diffusion process.
@@
To bridge the gap between leveraging 3D information for better video generation, we propose a method to augment and regularize video diffusion models using 3D point cloud. Specifically, our model enriches 2D videos with out-of-dimension data, which is absent in traditional RGB video models, without requiring full-scale 3D reconstruction. Instead of generating a complete mesh or point cloud for video diffusion model, we track the movement of 2D pixels in 3D space, creating a pseudo-3D representation of objects in the video. To enable such optimization, we further develop a 3D-aware video dataset, named PointVid, with segmented objects in pixel space and their corresponded 3D dynamic states. We fine-tune video diffusion models on the PointVid dataset to extend its generation capability to 3D shape and motion, enhancing the model to better understand object behavior in the physical world. Leveraging prior knowledge of 3D space, we utilize this additional information to guide pixel movements towards more physically plausible outcomes.
We demonstrate the effectiveness of our approach through extensive experiments, primarily focusing on handling complex scenarios with interacting objects and accurately resolving dynamic contacts. Guided by 3D trajectories, our model achieves superior quality in terms of smooth transitions of shape and motion compared to existing works, generating visually plausible videos. We summarize our contributions as follows:
  1. nosep
@@ 2 Related Work
Video Diffusion Models. Following the success of diffusion-based text-to-image (T2I) models [1, 27, 11, 22, 20, 24, 34, 21, 12, 17], many video generation models have been proposed in the literature [6, 32, 25, 14]. These models allow video generation following input conditioning such as text and image. Video diffusion models can be categorized into three categories. First, auto-regressive models like VideoPoet [19] follow the causal language model paradigm of next token prediction given previous tokens. These models encode the input conditioning into latent tokens using tokenizers such as MAGVIT [39].
@@
Second, pixel-space diffusion models [25, 14, 32] directly model the video synthesis on the pixel space to avoid artifacts arising in the compressed VAE latent spaces. While these models provide more photo-realism and better motion dynamics, they typically only offer low-resolution video generation due to a higher computational cost in generating high-resolution videos. These base models are augmented with super-resolution components to increase the video resolution resulting in deep cascaded diffusion models. Imagen-video [14], a UNet-based architecture, generates videos at a low-resolution and utilizes a sequence super-resolution models to increase the video resolution. Make-a-video [32] uses a T2I as prior to encode text input and trains a deep cascade of spatial and temporal layers to generate high resolution videos. SnapVideo [25] replaces the popular UNet architecture design with a FIT network for improved generation efficiency.
@@
Third, latent diffusion models transform the raw pixel space into a compressed spatio-temporal latent representation. It enables training a higher resolution and higher capacity base model instead of pixel-space diffusion models. These include Stable Video Diffusion [6], SORA [7], VideoCrafter2 [8], ModelScopeT2V [35], Mochi [33], CogVideoX [38, 15] etc. While latent models are typically single-stage pipelines, they can extend to deep cascade pipelines to offer even higher-quality video generations. MovieGen [28] proposed a transformer-based cascade latent diffusion model.
@@
While our approach is applicable to different types of video diffusion models, we mainly use ModelScopeT2V [35] as our base video-generation model and demonstrate how to incoporate 3D geometry and dynamics improve video generation.
@@
Dynamics Aware Video Generation. There have been various works [31, 23, 42] that instill physically plausible and temporally consistent dynamics in video generation. PhysGen [23] added simple control such as force or torque to an object in within the image, enabling simple physically plausible dynamics to the resulting video. MotionCraft [31] synthesized videos by incorporating optical flow in the noise diffusion process. This allows image animation with a temporally consistent video content. Physdreamer [42] tried distilling dynamic priors of static 3D objects from the video generation models. It creates a physical material field around the 3D object, thereby easily synthesizing 3D dynamics under arbritary forces. Although these works are able to incorporate simple dynamics to a static image, they assume that the underlying diffusion model understands and captures the object dynamics. This assumption is not necessarily true, as pointed by VideoPhy [4], current video generation models suffer from many inconsistencies, both physically implausible as well as temporally inconsistent. There have been other works along the line of incorporating camera intrinsics during video generation. 3D-Aware Video-Generation [2], first trains a model by jointing diffusing noise along a motion and content codes while sampling different camera poses. The model is trained using a generative adversarial network based discriminator which discriminates the resulting video, by difference between frames and time duration. This enables generating videos that are physically plausible.
@@
In contrast to the previous works, we incorporate 3D knowledge and dynamics by jointly diffusing the 3D geometry alongside the video modality. This helps improve the diffusion model understanding of the dynamics during training and improves the physical plausibility and temporal consistency of the generated videos.
@@ 3 Method
As mentioned earlier, current video models primarily learn moving objects as sequences of 2D pixels, without considering their true 3D states. Their understanding of the physical world is limited to a projected plane, making it challenging to resolve changes in shape and motion that are not visible in the projected view. To address this, we propose incorporating a partial 3D representation into the generation pipeline. The inclusion of 3D information as a second modality will not only enhance the generation capability of the video model but also guide the generation of RGB videos through joint optimization.
@@
Figure 3: PointVid Dataset Generation Workflow. Given an input video, we use the first frame as a reference frame and perform semantic segmentation to obtain masks for foreground objects. Next, we randomly sample pixels with a distribution favoring pixels inside foreground objects. We perform 3D point tracking on these queried pixels, and map these points to the input video frames. The resulting data point contains 3D coordinates of tracked foreground pixels while remaining pixels are zeroed out.
@@ 3.1 PointVid Dataset
We build our 3D-aware video dataset upon standard 3-channel videos of shape $[T,H,W,C]$, where $C=3$ encodes the RGB channels, height $H$ and width $W$ digitize 2D images into pixels, and $T$ represents the temporal dimension that sequences the pixel data over time.
The key challenge lies in selecting a 3D representation that is suitable for learning by a video model. Explicit 3D representations, such as meshes (surface or volumetric), point clouds, or voxels, can effectively model geometry but often have variable sizes due to irregular object shapes, complicating the diffusion training process. Moreover, aligning 3D points with 2D pixels in video remains a non-trivial task. In contrast, implicit representations like NeRF [26] typically require substantial time to convert into an explicit form for direct usage, making them impractical for training purposes.
Recent advancements in 3D point tracking [36] have made this task more feasible. Instead of representing moving objects as sequences of meshes or point clouds with arbitrary resolution, we focus exclusively on tracking the movements of pixels. Here, 2D pixels are unprojected into 3D space and tracked throughout the video. This allows the 3D locations of the pixels to be treated as three additional attributes, similar to color intensities. As a result, we can format the 3 spatial coordinates as extra channels, yielding a point tracking tensor $\mathcal{P}$ of shape $[T,H,W,C]$, matching the dimensions of the RGB video. In practice, we store the 3D tracking information using 2D pixel coordinates $(u,v)$ and depth $d$, requiring only a fixed camera unprojection to reconstruct the 3D world coordinates $(x,y,z)$. By using $(u,v,d)$, the first two channels correspond to the pixel’s position projected onto the 2D plane, which aligns precisely with the moving point in the RGB video, thus simplifying the task of learning the alignment between video and point tracking.
We visualize the workflow of generating point tracking in Figure 3. For each RGB video $\mathcal{V}$, we extract the first frame as a reference image and perform semantic segmentation using Grounded-SAM-2 [30, 29], yielding masks for the foreground objects. Next, we randomly sample pixels on the reference image with a main focus on those inside the foreground objects. These queried pixels are then tracked with trajectories of 3D Cartesian coordinates $(x,y,z)$ throughout the video using SpaTracker [36]. For reconstruction purposes, ZoeDepth [5] is employed for depth estimation. Finally, we postprocess the tracked results to align the shape with the video: we fill out the pixels in the foreground mask to store the pixel’s spatial positions, formatted as 3 channels, and discard pixels in the background (setting all background pixels to zero). Since we apply sparse queries for tracking efficiency, some foreground pixels may not have a corresponding trajectory; we use KDTree searching to interpolate the positions for such pixels.
@@
Table 1: Quantitative Evaluation. We evaluate various aspects of our method using the VBench [16] and VideoPhy [4] benchmarks. The evaluated metrics are as follows: (VBench) SC: subject consistency, BC: background consistency, MS: motion smoothness, AQ: aesthetic quality, IQ: imaging quality; (VideoPhy) PC: physical commonsense. By incorporating 3D knowledge, our video model shows substantial improvement in metrics such as physical commonsense, motion smoothness, and subject/background consistency. This demonstrates that our method generates significantly more temporally consistent and physically plausible videos.
I2VGen-XL 0.83247 0.89147 0.95706 0.44055 0.58532 0.32665
Ours 0.95892 0.95202 0.98456 0.43369 0.60423 0.37434
3.2 Point Augmentation

### v2

@@ Abstract
We present a novel video generation framework that integrates 3-dimensional geometry and dynamic awareness. To achieve this, we augment 2D videos with 3D point trajectories and align them in pixel space. The resulting 3D-aware video dataset, PointVid, is then used to fine-tune a latent diffusion model, enabling it to track 2D objects with 3D Cartesian coordinates. Building on this, we regularize the shape and motion of objects in the video to eliminate undesired artifacts, e.g., non-physical deformation. Consequently, we enhance the quality of generated RGB videos and alleviate common issues like object morphing, which are prevalent in current video models due to a lack of shape awareness. With our 3D augmentation and regularization, our model is capable of handling contact-rich scenarios such as task-oriented videos, where 3D information is essential for perceiving shape and motion of interacting solids. Our method can be seamlessly integrated into existing video diffusion models to improve their visual plausibility.
Figure 1: Comparison on Task-Oriented Videos. We present videos generated by different baselines (CogVideoX 1.5 yang2024cogvideox , SVD blattmann2023stablevideodiffusionscaling , DynamiCrafter xing2025dynamicrafter , LTX Video hacohen2024ltx and I2VGen-XL zhang2023i2vgen ) and compare them with our method. We use the same input conditions for all methods (except that SVD is conditioned only on the image). It can be observed that existing baselines often exhibit severe distortions of hands or objects during human hand-object interactions. In contrast, our method preserves the shapes of both the hand and object during such interactions.
@@ 1 Introduction
The recent development of video diffusion models has attracted significant research interests and seen tremendous progress blattmann2023stablevideodiffusionscaling ; singer2023makeavideo ; ho2022imagenvideohighdefinition . With vast amounts of training data and sophisticated architectures, video models have greatly improved in expressiveness and aesthetics videoworldsimulators2024 ; genmo2024mochi . Today, state-of-the-art video models can realistically represent content from the real world with high creativity, enabling various applications in the entertainment industry and scientific research Menapace_2024_CVPR . However, existing video diffusion models primarily focus on improving the appearance of content and the smoothness of motion. The models are trained in a way to understand the movements of 2-dimensional pixels, where the object movements in 3D space is approximated through the changing of RGB values.
@@
As a result, the underlying 3-dimensional shapes and motions, which represent how objects truly exist in the physical world, are not learned well in the existing video generation models. Even worse, we observe that when the video diffusion models neglect the underlying 3D information, they struggle to understand how objects occupy space, change shape and location, and interact with each other, which is essential for generating videos with complex motion. For instance, imaging a situation when perceiving a scene with a human performing specific tasks, e.g., hands tying shoelaces, viewpoint projection and occlusion make it impossible to capture the true 3D shape of the hand and the shoelaces using 2D pixels, complicating the modeling of how they make contact. Without understanding the underlying geometry, objects in the video may change shape arbitrarily or even appear and disappear suddenly, leading to the common issue of object morphing (as example results from prior arts in Figure 1). This problem is exacerbated for complex objects like humans; without 3D awareness, limbs and body parts may move unnaturally or change shape abruptly, resulting in dislocated arms or twisted bodies.
@@
To address the above issue, we aim to improve video diffusion models towards the better understanding of physical world, thus generating more reasonable shape and motion. Previous works have attempted to control shape and motion directly in 2D space savantaira2024motioncraft ; liu2024physgen . Nonetheless, without a comprehensive understanding of 3D geometry, these methods are limited to directing pixel movement confined within the 2D plane, making them unable to handle out-of-plane motions or fully resolve 3D dynamics liu2024physgen . One straightforward approach could be to model the complete 3D geometry of objects using representations such as NeRF mildenhall2021nerf or 3D Gaussian-Splatting kerbl20233d , essentially leading to 4D generation bahmani20244d ; zhao2023animate124 . While these methods provide accurate 3D understanding, they are computationally intensive and generally limited to generating simple scenes with only a few specific types of objects.
@@
To bridge the gap in leveraging 3D information for improved video generation, we propose a method to augment and regularize video diffusion models using 3D point clouds. Specifically, our model enriches 2D videos with out-of-plane data—absent in traditional RGB video models—without requiring full-scale 3D reconstruction. Instead of generating a complete mesh or point cloud for the video diffusion model, we track the movement of 2D pixels in 3D space, creating a pseudo-3D representation of objects in the video. To support this optimization, we develop a 3D-aware video dataset, named PointVid, which includes segmented objects in pixel space along with their corresponding 3D dynamic states. We fine-tune video diffusion models on the PointVid dataset to extend their generative capabilities to 3D shape and motion, enhancing the model’s ability to understand object behavior in the physical world. By leveraging prior knowledge of 3D space, we use this additional information to guide pixel movements toward more visually plausible outcomes.
Incorporating out-of-plane 3D information as a prior, our approach does not rely on specific architectures of the diffusion denoiser and thus can be integrated into various video models. To demonstrate this, we implement our method on two base models with distinct architectures: U-Net-based I2VGen-XL zhang2023i2vgen and DiT-based CogVideoX 1.5 yang2024cogvideox . Guided by 3D trajectories, our model achieves superior quality in terms of smooth transitions of shape and motion compared to existing works, generating visually plausible videos. We summarize our contributions as follows:
  1. 1.
@@ 2 Related Work
Video Diffusion Models. Following the success of diffusion-based text-to-image (T2I) models sd2.1 ; podell2023sdxl ; deepfloyd ; liu2024playgroundv3 ; flux ; liu2023hyperhuman ; kolors ; li2024snapfusion ; gao2024lumina ; kag2024ascan , many video generation models have been proposed in the literature blattmann2023stablevideodiffusionscaling ; singer2023makeavideo ; Menapace_2024_CVPR ; ho2022imagenvideohighdefinition . These models allow video generation following input conditioning such as text and image. Video diffusion models can be categorized into three categories. First, auto-regressive models like VideoPoet kondratyuk2024videopoetlargelanguagemodel follow the causal language model paradigm of next token prediction given previous tokens. These models encode the input conditioning into latent tokens using tokenizers such as MAGVIT yu2023magvit .
@@
Second, pixel-space diffusion models Menapace_2024_CVPR ; ho2022imagenvideohighdefinition ; singer2023makeavideo directly model the video synthesis on the pixel space to avoid artifacts arising in the compressed VAE latent spaces. While these models provide more photo-realism and better motion dynamics, they typically only offer low-resolution video generation due to a higher computational cost in generating high-resolution videos. These base models are augmented with super-resolution components to increase the video resolution resulting in deep cascaded diffusion models. Imagen-video ho2022imagenvideohighdefinition , a UNet-based architecture, generates videos at a low-resolution and utilizes a sequence super-resolution models to increase the video resolution. Make-a-video singer2023makeavideo uses a T2I as prior to encode text input and trains a deep cascade of spatial and temporal layers to generate high resolution videos. SnapVideo Menapace_2024_CVPR replaces the popular UNet architecture design with a FIT network for improved generation efficiency.
@@
Third, latent diffusion models transform the raw pixel space into a compressed spatio-temporal latent representation. It enables training a higher resolution and higher capacity base model instead of pixel-space diffusion models. These include Stable Video Diffusion blattmann2023stablevideodiffusionscaling , SORA videoworldsimulators2024 , VideoCrafter2 chen2024videocrafter2 , ModelScopeT2V wang2023modelscope , Mochi genmo2024mochi , LTX Video hacohen2024ltx , CogVideoX yang2024cogvideox ; hong2022cogvideo etc. While latent models are typically single-stage pipelines, they can extend to deep cascade pipelines to offer even higher-quality video generations. MovieGen polyak2024moviegencastmedia proposed a transformer-based cascade latent diffusion model.
@@
While our approach is applicable to various types of video diffusion models, we select two models with distinct underlying architectures— I2VGen-XL zhang2023i2vgen and CogVideoX-1.5 yang2024cogvideox —as our base models, and demonstrate how incorporating 3D geometry and dynamics improves video generation.
@@
Dynamics Aware Video Generation. There have been various works savantaira2024motioncraft ; liu2024physgen ; zhang2024physdreamer that instill physically plausible and temporally consistent dynamics in video generation. PhysGen liu2024physgen added simple control such as force or torque to an object in within the image, enabling simple physically plausible dynamics to the resulting video. MotionCraft savantaira2024motioncraft synthesized videos by incorporating optical flow in the noise diffusion process. This allows image animation with a temporally consistent video content. Physdreamer zhang2024physdreamer tried distilling dynamic priors of static 3D objects from the video generation models. It creates a physical material field around the 3D object, thereby easily synthesizing 3D dynamics under arbritary forces. Although these works are able to incorporate simple dynamics to a static image, they assume that the underlying diffusion model understands and captures the object dynamics. This assumption is not necessarily true, as pointed by VideoPhy bansal2024videophy , current video generation models suffer from many inconsistencies, both physically implausible as well as temporally inconsistent. There have been other works along the line of incorporating camera intrinsics during video generation. 3D-Aware Video-Generation bahmani20233dawarevideogeneration , first trains a model by jointing diffusing noise along a motion and content codes while sampling different camera poses. The model is trained using a generative adversarial network based discriminator which discriminates the resulting video, by difference between frames and time duration.
@@
In contrast to previous works, we incorporate 3D knowledge by jointly diffusing point clouds alongside the video modality. This enhances the diffusion model’s understanding of 3D world during training and improves the visual plausibility and spatio-temporal consistency of the generated videos.
Figure 2: Method Overview. During training, we sample video-point pairs, concatenate them along the channel dimensions, and use the augmented data to train a latent diffusion model. We introduce cross-attention between video and point data in corresponding channels to enhance alignment between the two modalities. The model predicts both RGB video and 3D points, leveraging the 3D information to further regularize video generation by applying a misalignment penalty during the diffusion process. During inference, the model generates both video and points from random noise, conditioned on a text-image prompt.
@@ 3 Method
As mentioned earlier, current video models primarily learn moving objects as sequences of 2D pixels, without considering their true 3D states. Their understanding of the physical world is limited to a projected plane, making it challenging to resolve changes in shape and motion that are not visible in the projected view. To address this, we propose incorporating a partial 3D representation into the generation pipeline. The inclusion of 3D information as a second modality will not only enhance the capability of the video model but also guide the generation of RGB videos through joint optimization.
@@ 3.1 PointVid Dataset
We build our 3D-aware PointVid dataset upon standard 3-channel videos of shape $[T,H,W,C]$, where $C=3$ encodes the RGB channels, height $H$ and width $W$ digitize 2D images into pixels, and $T$ represents the temporal dimension that sequences the pixel data over time.
The key challenge lies in selecting a 3D representation that is suitable for learning by a video model. Explicit 3D representations, such as meshes (surface or volumetric), point clouds, or voxels, can effectively model geometry but often have variable sizes due to irregular object shapes, complicating the diffusion training process. Moreover, aligning 3D points with 2D pixels in video remains a non-trivial task. In contrast, implicit representations like NeRF mildenhall2021nerf typically require substantial time to convert into an explicit form for direct usage, making them impractical for training purposes.
Recent advancements in 3D point tracking xiao2024spatialtracker have made this task more feasible. Instead of representing moving objects as sequences of meshes or point clouds with arbitrary resolution, we focus exclusively on tracking the movements of pixels. Here, 2D pixels are unprojected into 3D space and tracked throughout the video. This allows the 3D locations of the pixels to be treated as three additional attributes, similar to color intensities. As a result, we can format the 3 spatial coordinates as extra channels, yielding a point tracking tensor $\mathcal{P}$ of shape $[T,H,W,C]$, matching the dimensions of the RGB video. In practice, we store the 3D tracking information using 2D pixel coordinates $(u,v)$ and depth $d$, requiring only a simple camera unprojection to reconstruct the 3D world coordinates $(x,y,z)$. By using $(u,v,d)$, the first two channels correspond to the pixel’s position projected onto the 2D plane, which aligns precisely with the moving point in the RGB video, thus simplifying the task of learning the alignment between video and point tracking.
We visualize the workflow of generating point tracking in Figure 3. For each RGB video $\mathcal{V}$, we extract the first frame as a reference image and perform semantic segmentation using Grounded-SAM-2 ren2024grounded ; ravi2024sam2segmentimages , yielding masks for the foreground objects. Next, we randomly sample pixels on the reference image with a main focus on those inside the foreground objects. These queried pixels are then tracked with trajectories of 3D Cartesian coordinates $(x,y,z)$ throughout the video using SpaTracker xiao2024spatialtracker . Finally, we post-process the tracked results to align the shape with the video: we fill out the pixels in the foreground mask to store the pixel’s spatial positions, formatted as 3 channels, and discard pixels in the background (setting all background pixels to zero). Since we apply sparse queries for tracking efficiency, some foreground pixels may not have a corresponding trajectory; we use KDTree searching to interpolate the positions for such pixels.
@@
Figure 3: PointVid Dataset Generation. Given an input video, we use the first frame as a reference frame and perform semantic segmentation to obtain masks for foreground objects. Next, we randomly sample pixels with a distribution favoring pixels inside foreground objects. We perform 3D point tracking on these queried pixels, and map these points to the input video frames. The resulting data point contains 3D coordinates of tracked foreground pixels while remaining pixels are zeroed out.
3.2 Point Augmentation in Diffusion Denoiser